Give the following frequency distribution with some missing frequencies

Class

10-20

20-30

30-40

40-50

50-60

60-70

70-80


Frequency

185

f2

34

180

136

f6

50



If the total frequency is 686 and the midian is 42.6 Find out the missing frequencies f2 and f6.


1
Expert's answer
2021-10-21T11:59:40-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 Class & Frequency & C.F. \\ \hline
 10-20 & 185 & 185 \\
 \hdashline
 20-30 & f2 & 185+f2 \\
 \hdashline
 30-40 & 34 & 219+f2 \\
 \hdashline
 40-50 & 180& 399+f2 \\
 \hdashline
 50-60 & 136& 535+f2 \\
 \hdashline
 60-70 & f6 & 535+f2+f6 \\
 \hdashline
 70-80 & 50 & 585+f2+f6 \\
 \hdashline
\end{array}"

The total frequency is 686


"585+f2+f6=686"

"f2+f6=101"

"median=42.6" Then median class is "40-50."

"l=40, h=10, f=180," and "C=219+f2"


"median=l+\dfrac{N/2-C}{f}\times h"

"42.6=40+\dfrac{686/2-219-f2}{180}\times 10"

"124-f2=46.8"

"f2=77"

"f6=101-77=24"